OpenStudy (surryyy):

We toss two dice 1,000 times. How many times do we expect to have the sum of the two dice equal to 4? Answer about 250 about 167 about 83 about 42

ganeshie8 (ganeshie8):

1 + 3 = 4 2 + 2 = 4 3 + 1 = 4

ganeshie8 (ganeshie8):

looks there are 3 ways to get a 4 ?

ganeshie8 (ganeshie8):

number of favorable outcomes = 3

ganeshie8 (ganeshie8):

whats the total number of outcomes when you roll 2 dice ?